OpenStudy (anonymous):

Find an equation of the given line containing the given pair of points . (-8,-8) and (6,6) how do you do these problems

OpenStudy (aroub):

ok first u have to find the slope.. do u know how?

OpenStudy (anonymous):

not really

OpenStudy (aroub):

ok here is the formula : (x1,y1) , (x2,y2) y2-y1/ x2-x1

OpenStudy (aroub):

now try using the points u have..

OpenStudy (aroub):

trying??

OpenStudy (anonymous):

so it will be -8-6 then 6-8 right

OpenStudy (aroub):

no..

OpenStudy (aroub):

and it will be the same for both anyway.. b/c u have same points

OpenStudy (aroub):

ok see.. u have (-8,-8) (6,6) , just apply the formula u will get 6+8/6+8 14/14 1 do u get it? or not really?

OpenStudy (anonymous):

do u need help

OpenStudy (anonymous):

so im adding 6+8 which is 14 so the answer would be 14

OpenStudy (aroub):

yes its 14 but 14/14 is 1 right? do u agree??

OpenStudy (anonymous):

Required eq. of line is: (y+8)/(6+8)=(x+8)/(6+8) =>y+8=x+8 =>y=x

OpenStudy (aroub):

ok u have a slope of 1 , now we want to find the line, i prefer using y=mx=b (or if u want u can use y-y1=m(x-x1) but now now we will use y=mx+b ok? y=mx+b ( u can choose any point u want from the 2 points) ill choose (6,6) (x,y) 6=1(6)+b 6=6+b b=0 (we found the y-intercept which is zero ao the equation is y=x